# Unifying the Three Algebraic Approaches to the CSP via Minimal Taylor Algebras

### v5

@@ 9.4 Absorption and edges
If the algebra $\mathbf{A}$ is minimal Taylor we can stop the proof here, because we have already obtained the subpower $B(x_{1})\vee B(x_{2})\vee B(x_{3})$ (see Theorem 5.7). For general Taylor algebras, we can by induction obtain the subpower $B(x_{1})\vee B(x_{2})\vee\dots\vee B(x_{n})$ for every $n$, in a completely analogous way. ∎
@@
Proof.
That absorbing sets are stable under semilattice and abelian edges and that (a) implies (b) follow from Lemma 9.8. That (b) implies (a) follows from Theorem 9.13. Indeed, if $\{b\}$ does not absorb $\mathbf{A}$ then we take a minimal subalgebra $\mathbf{B}$ of $\mathbf{A}$ such that $b\in B$ and $\{b\}$ does not absorb $\mathbf{B}$, and apply the theorem. Either $\mathbf{B}$ has a nontrivial abelian quotient, in which case $\{b\}$ is not stable under abelian edges, or $\mathbf{B}$ has a nonempty projective subuniverse $C$ that does not contain $b$. In the latter case $C\trianglelefteq_{2}\mathbf{A}$ by Proposition 8.2 and there exists a (minimal) semilattice edge $(b,c)$ with $c\in C$ by Proposition 9.7, therefore $\{b\}$ is not stable under semilattice edges. ∎
@@
Let $t$ be an $n$-ary cyclic term operation of a minimal Taylor algebra $\mathbf{A}$ and $B\trianglelefteq_{3}\mathbf{A}$. Then for any $m\geq n/2$ and any ${\mathbf{a}}\in A^{n}$ such that $a_{1},\dots,a_{m}\in B$ we have $t({\mathbf{a}})\in B$.
Proof.
Let ${\mathbf{b}}$ be the tuple obtained by cyclically shifting the tuple ${\mathbf{a}}$ by $n-m$ positions. Since $m\geq n/2$, each pair $(a_{i},b_{i})$ is in the relation $B(x_{1})\vee B(x_{2})$; and since $t$ is cyclic, we have $t({\mathbf{a}})=t({\mathbf{b}})$. But $t$ is compatible with $B(x_{1})\vee B(x_{2})$ by Theorem 5.10, therefore $t({\mathbf{a}})\in B$. ∎
@@
Proof.
Choose positive integers $n,k,l$ such that $n\equiv 1\pmod{|A|!}$, $k\equiv 1\pmod{|A|!}$, $2k+l=n$, and $2k\geq l$. Let $m=k$. Note that $\mathbf{A}$ has a cyclic operation $t$ of arity $n$: every prime divisor $p$ of $n$ is greater than $|A|$ and thus there exists a cyclic term operation $t_{p}$ of arity $p$ by Theorem 3.5. The operation $t$ can then be obtained by a star composition of the $t_{p}$s.
@@
Because $\mathbf{E}/\theta$ and $\mathbf{A}$ are both minimal Taylor algebras, the operation $f$ satisfies the second and the fourth item by using the cyclicity of $t$ and Proposition 9.14, and it satisfies the first and the fifth item by Theorem 5.7. Note that for these claims we actually only need $k+l,k+m,l+m\geq n/2$.
For the third item recall item (c) in Theorem 5.13 and note that a cyclic operation in an affine Mal’cev algebra of $\mathbb{Z}/{q}$ is equal to $\sum_{i=1}^{n}ax_{i}\pmod{q}$ where $na\equiv 1\pmod{q}$. By simple arithmetic we conclude that if the conditions on $k,l,m$ hold, then the ternary operation $f$ is $x-y+z$. Indeed, since $q$ divides $|A|!$, we get $k=m\equiv 1\equiv n\pmod{q}$, $a\equiv na\equiv 1\pmod{q}$, and $l=n-2k\equiv-1\pmod{q}$, so $f(x,y,z)\equiv akx+aly+azm\equiv x-y+z\pmod{q}$. ∎
@@
Let $\mathbf{A}$ be a minimal Taylor algebra generated by $a,b\in A$. Suppose that $C\trianglelefteq_{3}\mathbf{A}$ is nontrivial and 2-absorbs a subalgebra of $\mathbf{A}$ containing $a$ and a subalgebra of $\mathbf{A}$ containing $b$. Then $\mathbf{A}$ has a proper 3-absorbing subuniverse containing $a$ or $b$.
Proof.
For simplicity, we will say that a set 2-absorbs $a$ (or $b$) if it 2-absorbs a subalgebra containing $a$ (or $b$). Let $S=\operatorname{Sg}_{\mathbf{A}^{2}}((a,b),(b,a))$. As a first step we obtain a nontrivial 3-absorbing subuniverse $D$ that 2-absorbs $a$ and $b$ and such that $S\cap(C\times D)$ is nonempty: take $D=C$ if $C+S=A$, or $D=C+S$ otherwise. Fix $(c,d)\in S\cap(C\times D)$ and let $f$ be a witness for $(c,d)\in S$, that is, $f(a,b)=c$ and $f(b,a)=d$. Next observe that if $a$ is in $C$ or $D$, then the goal is reached. Suppose henceforth that $C\cup D$ does not contain $a$.
By Theorem 9.5, $f\rightsquigarrow_{C}f^{\prime}$ and $f\rightsquigarrow_{D}f^{\prime}$ for some monotone selfdual binary operation, i.e., $f^{\prime}$ is one of the two projections. Suppose $f^{\prime}$ is the first projection, so $f(C,A)\subseteq C$ and $f(D,A)\subseteq D$. From these inclusions, from $f(a,b),f(b,a)\in C\cup D$, and from $f(\{a,b\},C)\subseteq C$, $f(\{a,b\},D)\subseteq D$ (by strong projectivity from Theorem 5.7) it follows that $C\cup D\cup\{b\}$ absorbs the set $E=C\cup D\cup\{a,b\}$ by $f$. It is therefore enough to show that $E$ is a subuniverse of $\mathbf{A}$ – then $E=\operatorname{Sg}(a,b)$ and the proof will be concluded.
@@
and let $h$ be the cyclic composition of $t$ and $s$. The result of applying $s$ to a tuple ${\mathbf{e}}\in E^{p}$ is, by the properties above, in $C\cup D$ whenever ${\mathbf{e}}$ is not the constant tuple of $a$’s or $b$’s. We then have $h({\mathbf{e}})\in t(C\cup D,\dots,C\cup D)$, which is by item (1) of Proposition 5.12 a subset of $C\cup D$. Since $h$ is a cyclic operation and it preserves $E$, we see that $E$ is a subuniverse of $\mathbf{A}$ by Proposition 5.3, which concludes the proof. ∎
@@
First we will prove that every reflexive relation $S\leq_{sd}\mathbf{A}^{2}$ is either the equality or the full relation. Suppose not. Then, by simplicity of $\mathbf{A}$, $S$ is linked. By replacing $S$ by $S-S$, perhaps multiple times if necessary, we can further assume that $S-S=A^{2}$ while $S$ is still not full. Then there is some $c\in A$ such that $(a,c),(b,c)\in S$, and since $A=\operatorname{Sg}(a,b)$, we see that $S$ has a nontrivial right center $C$ and is still reflexive.
@@
We claim that $C$ 2-absorbs the subalgebra $\mathbf{D}$ of $\mathbf{A}$ with universe $a+S$ (which contains $a$ since $S$ is reflexive). By Proposition 9.10. we just need to check that for any $d\in D\setminus C$ and any $c\in C$, $\operatorname{Sg}(d,c)$ has a proper 2-absorbing subuniverse. To see this, note that the binary relation $T=S\cap(A\times\operatorname{Sg}(d,c))$ is a subdirect subuniverse of $\mathbf{A}\times\operatorname{Sg}(d,c)$ (since $c$ is in the right center of $S$) and $a$ is contained in the left center of $T$ (since $d,c\in a+S$). Since $d\not\in C$, we see that $T$ is a proper subset of $A\times\operatorname{Sg}(d,c)$, so the left center of $T$ is a proper subalgebra of $\mathbf{A}$. Thus if $\operatorname{Sg}(d,c)$ has no proper 2-absorbing subuniverse, then the left center of $T$ is a center of $\mathbf{A}$, and $a$ is contained in a proper 3-absorbing subuniverse of $\mathbf{A}$ by Proposition 4.3, a contradiction. The same argument shows that $C$ 2-absorbs a subalgebra containing $b$. Application of Lemma 9.15 then gives us a contradiction, so we have proved that every reflexive subdirect subuniverse of $\mathbf{A}^{2}$ is the equality or the full relation.
@@
Observe next that $S=\operatorname{Sg}_{\mathbf{A}^{2}}((a,b),(b,a))$ is a graph of a bijection. Indeed, it cannot be full as otherwise $(b,b)\in S$ implies that $(a,b)$ is an s-edge and then $\{b\}$ 2-absorbs $A=\{a,b\}$. If it is proper and linked, then we can use the same argument as above with the little tweak that $D$ may contain $b$ (not $a$) since $S$ is not reflexive (but contains $(a,b)$).
@@
Finally, consider the subdirect symmetric ternary relation $R=\operatorname{Sg}_{\mathbf{A}^{3}}\{(a,a,b),(a,b,a),(b,a,a)\}.$ Since the projection $\operatorname{proj}_{12}(R)$ of $R$ onto the first two coordinates contains $(a,a),(a,b)$ and $\operatorname{Sg}(a,b)=A$, we see that $\operatorname{proj}_{12}(R)$ has a left center that contains $a$. The center cannot be proper since otherwise we get a nontrivial 2-absorbing subuniverse or the center of $\operatorname{proj}_{12}(R)$ is a Taylor center of $\mathbf{A}$ and we apply Proposition 4.3 to get a contradiction. Therefore $\operatorname{proj}_{12}(R)=A^{2}$. If $(a,a,a)\in R$, then the same argument as in Lemma 9.11 shows that $(a,b)$ is a majority edge, which gives us a proper 3-absorbing subuniverse containing $a$ (and also one containing $b$), a contradiction. The binary relation $T(x,y)\equiv R(a,x,y)$ is thus subdirect (as the projection of $R$ onto any two coordinates is full) but not full (as $(a,a)\not\in T$). Observe that $T+S$ is reflexive but cannot be full (as otherwise $T=(T+S)+S$ is full as well), therefore it is the equality relation and hence $T$ is a graph of a bijection. It follows that the strongly projective subuniverse $B$ from Lemma 9.12 is neither empty (as $\mathbf{A}$ is not abelian) nor equal to $A$ (as $a\not\in B$). Theorem 5.7 shows that $B$ 2-absorbs $\mathbf{A}$, which is a contradiction. ∎
@@
Let $\mathbf{A}$ be a minimal Taylor algebra and $f_{1},...,f_{n}$ any collection of term operations of $\mathbf{A}$ such that for each edge $(a,b)$ with witnessing maximal congruence $\theta$ of $\operatorname{Sg}_{\mathbf{A}}(a,b)$, at least one $f_{i}$ acts nontrivially on $\operatorname{Sg}_{\mathbf{A}}(a,b)/\theta$ (that is, does not act as a projection). Then $f_{1},...,f_{n}$ generate the clone of $\mathbf{A}$.
Proof.
Let $\mathbf{G}$ be the reduct of $\mathbf{A}$ with basic operations $f_{1},...,f_{n}$. By Proposition 2.2 it is enough to show that for any $a,b\in G$, there is no two-element quotient $\operatorname{Sg}_{\mathbf{G}}(a,b)/\alpha$ such that each $f_{i}$ acts as a projection. Suppose for contradiction that there was such a pair $a,b$ and congruence $\alpha$ on $\operatorname{Sg}_{\mathbf{G}}(a,b)$, and choose $a,b$ such that $\operatorname{Sg}_{\mathbf{G}}(a,b)$ is inclusion minimal. Note that $a/\alpha$, $b/\alpha$, and $\operatorname{Sg}_{\mathbf{G}}(a,b)$ might not be subuniverses of $\mathbf{A}$.
@@
If there is a congruence $\theta$ such that $\operatorname{Sg}_{\mathbf{A}}(a,b)/\theta$ is affine of prime order, then by assumption some $f_{i}$ acts nontrivially on $\operatorname{Sg}_{\mathbf{A}}(a,b)/\theta$, so there is some ternary term operation $p\in\operatorname{Clo}_{3}(\mathbf{G})$ such that the restriction of $p$ to $\operatorname{Sg}_{\mathbf{A}}(a,b)/\theta$ is Mal’cev (by the fact that the affine Mal’cev algebra of $\mathbb{Z}/{p}$ has minimal clone which follows e.g. from the description of term operations of this algebra). Since $f_{i}$ acts as a projection on $\operatorname{Sg}_{\mathbf{G}}(a,b)/\alpha$, $p$ also acts as a projection on $\operatorname{Sg}_{\mathbf{G}}(a,b)/\alpha$. Suppose without loss of generality that $p$ does not act like third projection on $\operatorname{Sg}_{\mathbf{G}}(a,b)/\alpha$. Then we have
@@
But then $\operatorname{Sg}_{\mathbf{G}}(p(a,a,b),b)/\alpha=\operatorname{Sg}_{\mathbf{G}}(a,b)/\alpha$ and
$\operatorname{Sg}_{\mathbf{G}}(p(a,a,b),b)\subseteq\operatorname{Sg}_{\mathbf{G}}(a,b)\cap b/\theta,$
contradicting the minimality of $\operatorname{Sg}_{\mathbf{G}}(a,b)$ (as the latter set does not contain $a$ because $\theta$ separates $a$ and $b$).
Now suppose that $a$ is contained in a proper 3-absorbing subalgebra $\mathbf{C}$ of $\operatorname{Sg}_{\mathbf{A}}(a,b)$. We show that this absorption is witnessed by an operation in $\operatorname{Clo}(\mathbf{G})$.
If $C$ is a 2-absorbing subuniverse of $\operatorname{Sg}_{\mathbf{A}}(a,b)$, then there is some s-edge going into $C$ (by Proposition 9.7), and then some $f_{i}$ acts nontrivially on this edge, so there is a binary term operation $s$ of $\mathbf{G}$ which acts as a semilattice operation on this edge. Since both coordinates of $s$ are essential and $C$ is a strongly projective subuniverse of $\operatorname{Sg}_{\mathbf{A}}(a,b)$ by Theorem 5.7, this operation witnesses the 2-absorption $C\trianglelefteq_{2}\operatorname{Sg}_{\mathbf{A}}(a,b)$.
If $C$ is not a 2-absorbing subuniverse of $\operatorname{Sg}_{\mathbf{A}}(a,b)$, then $C$ is not stable under majority edges by Lemma 9.8 and Theorem 5.19. Let $(c,d)$ with witnessing congruence $\beta$ be a majority edge such that $c\in C$ and $d/\beta\cap C=\emptyset$. Some $f_{i}$ acts nontrivially on $\operatorname{Sg}_{\mathbf{A}}(c,d)/\beta$, so there is a ternary term operation $m$ of $\mathbf{G}$ that acts as the majority operation on $\operatorname{Sg}_{\mathbf{A}}(c,d)/\beta$. We claim that $m$ witnesses $C\trianglelefteq_{3}\operatorname{Sg}_{\mathbf{A}}(a,b)$. Indeed, $m\rightsquigarrow_{C}g$ for some $g\in\operatorname{Clo}(\{0,1\},\operatorname{maj})$ by Theorem 5.10, but $g$ cannot be a projection: e.g., if $g$ is the first projection, then $m(c,d,d)\in C$ (because $m$ is $\rightsquigarrow_{C}$-related to the first projection) and $m(c,d,d)\in d/\beta$ (because $m$ acts as the majority operation on the majority edge), a contradiction. Therefore $g=\operatorname{maj}$ and $m\rightsquigarrow_{C}\operatorname{maj}$ implies the claim.
In both cases, we see that there is some ternary operation $t\in\operatorname{Clo}(\mathbf{G})$ which witnesses the 3-absorption $C\trianglelefteq_{3}\operatorname{Sg}_{\mathbf{A}}(a,b)$, and we may suppose without loss of generality that $t$ acts like the first projection on $\operatorname{Sg}_{\mathbf{G}}(a,b)/\alpha$. Then we have
@@
But then $\operatorname{Sg}_{\mathbf{G}}(a,t(b,a,a))/\alpha=\operatorname{Sg}_{\mathbf{G}}(a,b)/\alpha$ and
$\operatorname{Sg}_{\mathbf{G}}(a,t(b,a,a))\subseteq\operatorname{Sg}_{\mathbf{G}}(a,b)\cap C,$
contradicting the minimality of $\operatorname{Sg}_{\mathbf{G}}(a,b)$ again, and concluding the proof. ∎
@@
Proof.
This is a consequence of Theorem 9.18. ∎

### v6

@@ 9.4 Absorption and edges
If the algebra is minimal Taylor we can stop the proof here, because we have already obtained the subpower $B(x_{1})\vee B(x_{2})\vee B(x_{3})$ (see Theorem 5.7). For general Taylor algebras, we can by induction obtain the subpower $B(x_{1})\vee B(x_{2})\vee\dots\vee B(x_{n})$ for every $n$, in a completely analogous way.
@@
Proof 9.37.
That absorbing sets are stable under semilattice and abelian edges and that (a) implies (b) follow from Lemma 9.25. That (b) implies (a) follows from Theorem 9.35. Indeed, if $\{b\}$ does not absorb then we take a minimal subalgebra of such that $b\in B$ and $\{b\}$ does not absorb , and apply the theorem. Either has a nontrivial abelian quotient, in which case $\{b\}$ is not stable under abelian edges, or has a nonempty projective subuniverse $C$ that does not contain $b$. In the latter case $C\trianglelefteq_{2}$ by Proposition 8.2 and there exists a (minimal) semilattice edge $(b,c)$ with $c\in C$ by Proposition 9.22, therefore $\{b\}$ is not stable under semilattice edges.
@@
Let $t$ be an $n$-ary cyclic term operation of a minimal Taylor algebra and $B\trianglelefteq_{3}$. Then for any $m\geq n/2$ and any ${\mathbf{a}}\in A^{n}$ such that $a_{1},\dots,a_{m}\in B$ we have $t({\mathbf{a}})\in B$.
Proof 9.39.
Let ${\mathbf{b}}$ be the tuple obtained by cyclically shifting the tuple ${\mathbf{a}}$ by $n-m$ positions. Since $m\geq n/2$, each pair $(a_{i},b_{i})$ is in the relation $B(x_{1})\vee B(x_{2})$; and since $t$ is cyclic, we have $t({\mathbf{a}})=t({\mathbf{b}})$. But $t$ is compatible with $B(x_{1})\vee B(x_{2})$ by Theorem 5.10, therefore $t({\mathbf{a}})\in B$.
@@
Proof 9.40.
Choose positive integers $n,k,l$ such that $n\equiv 1\pmod{|A|!}$, $k\equiv 1\pmod{|A|!}$, $2k+l=n$, and $2k\geq l$. Let $m=k$. Note that has a cyclic operation $t$ of arity $n$: every prime divisor $p$ of $n$ is greater than $|A|$ and thus there exists a cyclic term operation $t_{p}$ of arity $p$ by Theorem 3.5. The operation $t$ can then be obtained by a star composition of the $t_{p}$s.
@@
Because and are both minimal Taylor algebras, the operation $f$ satisfies the second and the fourth item by using the cyclicity of $t$ and Proposition 9.38, and it satisfies the first and the fifth item by Theorem 5.7. Note that for these claims we actually only need $k+l,k+m,l+m\geq n/2$.
For the third item recall item (c) in Theorem 5.13 and note that a cyclic operation in an affine Mal’cev algebra of $\mathbb{Z}/{q}$ is equal to $\sum_{i=1}^{n}ax_{i}\pmod{q}$ where $na\equiv 1\pmod{q}$. By simple arithmetic we conclude that if the conditions on $k,l,m$ hold, then the ternary operation $f$ is $x-y+z$. Indeed, since $q$ divides $|A|!$, we get $k=m\equiv 1\equiv n\pmod{q}$, $a\equiv na\equiv 1\pmod{q}$, and $l=n-2k\equiv-1\pmod{q}$, so $f(x,y,z)\equiv akx+aly+azm\equiv x-y+z\pmod{q}$.
@@
Let be a minimal Taylor algebra generated by $a,b\in A$. Suppose that $C\trianglelefteq_{3}$ is nontrivial and 2-absorbs a subalgebra of containing $a$ and a subalgebra of containing $b$. Then has a proper 3-absorbing subuniverse containing $a$ or $b$.
Proof 9.42.
For simplicity, we will say that a set 2-absorbs $a$ (or $b$) if it 2-absorbs a subalgebra containing $a$ (or $b$). Let $S=\operatorname{Sg}((a,b),(b,a))$. As a first step we obtain a nontrivial 3-absorbing subuniverse $D$ that 2-absorbs $a$ and $b$ and such that $S\cap(C\times D)$ is nonempty: take $D=C$ if $C+S=A$, or $D=C+S$ otherwise. Fix $(c,d)\in S\cap(C\times D)$ and let $f$ be a witness for $(c,d)\in S$, that is, $f(a,b)=c$ and $f(b,a)=d$. Next observe that if $a$ is in $C$ or $D$, then the goal is reached. Suppose henceforth that $C\cup D$ does not contain $a$.
By Theorem 9.14, $f\rightsquigarrow_{C}f^{\prime}$ and $f\rightsquigarrow_{D}f^{\prime}$ for some monotone selfdual binary operation, i.e., $f^{\prime}$ is one of the two projections. Suppose $f^{\prime}$ is the first projection, so $f(C,A)\subseteq C$ and $f(D,A)\subseteq D$. From these inclusions, from $f(a,b),f(b,a)\in C\cup D$, and from $f(\{a,b\},C)\subseteq C$, $f(\{a,b\},D)\subseteq D$ (by strong projectivity from Theorem 5.7) it follows that $C\cup D\cup\{b\}$ absorbs the set $E=C\cup D\cup\{a,b\}$ by $f$. It is therefore enough to show that $E$ is a subuniverse of – then $E=\operatorname{Sg}(a,b)$ and the proof will be concluded.
@@
and let $h$ be the cyclic composition of $t$ and $s$. The result of applying $s$ to a tuple ${\mathbf{e}}\in E^{p}$ is, by the properties above, in $C\cup D$ whenever ${\mathbf{e}}$ is not the constant tuple of $a$’s or $b$’s. We then have $h({\mathbf{e}})\in t(C\cup D,\dots,C\cup D)$, which is by item (1) of Proposition 5.12 a subset of $C\cup D$. Since $h$ is a cyclic operation and it preserves $E$, we see that $E$ is a subuniverse of by Proposition 5.3, which concludes the proof.
@@
First we will prove that every reflexive relation $S\leq_{sd}$ is either the equality or the full relation. Suppose not. Then, by simplicity of , $S$ is linked. By replacing $S$ by $S-S$, perhaps multiple times if necessary, we can further assume that $S-S=A^{2}$ while $S$ is still not full. Then there is some $c\in A$ such that $(a,c),(b,c)\in S$, and since $A=\operatorname{Sg}(a,b)$, we see that $S$ has a nontrivial right center $C$ and is still reflexive.
@@
We claim that $C$ 2-absorbs the subalgebra of with universe $a+S$ (which contains $a$ since $S$ is reflexive). By Proposition 9.28. we just need to check that for any $d\in D\setminus C$ and any $c\in C$, $\operatorname{Sg}(d,c)$ has a proper 2-absorbing subuniverse. To see this, note that the binary relation $T=S\cap(A\times\operatorname{Sg}(d,c))$ is a subdirect subuniverse of $\mathbf{A}\times\operatorname{Sg}(d,c)$ (since $c$ is in the right center of $S$) and $a$ is contained in the left center of $T$ (since $d,c\in a+S$). Since $d\not\in C$, we see that $T$ is a proper subset of $A\times\operatorname{Sg}(d,c)$, so the left center of $T$ is a proper subalgebra of $\mathbf{A}$. Thus if $\operatorname{Sg}(d,c)$ has no proper 2-absorbing subuniverse, then the left center of $T$ is a center of , and $a$ is contained in a proper 3-absorbing subuniverse of $\mathbf{A}$ by Proposition 4.3, a contradiction. The same argument shows that $C$ 2-absorbs a subalgebra containing $b$. Application of Lemma 9.41 then gives us a contradiction, so we have proved that every reflexive subdirect subuniverse of  is the equality or the full relation.
@@
Observe next that $S=\operatorname{Sg}((a,b),(b,a))$ is a graph of a bijection. Indeed, it cannot be full as otherwise $(b,b)\in S$ implies that $(a,b)$ is an s-edge and then $\{b\}$ 2-absorbs $A=\{a,b\}$. If it is proper and linked, then we can use the same argument as above with the little tweak that $D$ may contain $b$ (not $a$) since $S$ is not reflexive (but contains $(a,b)$).
@@
Finally, consider the subdirect symmetric relation $R=\operatorname{Sg}\{(a,a,b),(a,b,a),(b,a,a)\}.$ Since the projection $\operatorname{proj}_{12}(R)$ of $R$ onto the first two coordinates contains $(a,a),(a,b)$ and $\operatorname{Sg}(a,b)=A$, we see that $\operatorname{proj}_{12}(R)$ has a left center that contains $a$. The center cannot be proper since otherwise we get a nontrivial 2-absorbing subuniverse or the center of $\operatorname{proj}_{12}(R)$ is a Taylor center of and we apply Proposition 4.3 to get a contradiction. Therefore $\operatorname{proj}_{12}(R)=A^{2}$. If $(a,a,a)\in R$, then the same argument as in Lemma 9.30 shows that $(a,b)$ is a majority edge, which gives us a proper 3-absorbing subuniverse containing $a$ (and also one containing $b$), a contradiction. The binary relation $T(x,y)\equiv R(a,x,y)$ is thus subdirect (as the projection of $R$ onto any two coordinates is full) but not full (as $(a,a)\not\in T$). Observe that $T+S$ is reflexive but cannot be full (as otherwise $T=(T+S)+S$ is full as well), therefore it is the equality relation and hence $T$ is a graph of a bijection. It follows that the strongly projective subuniverse $B$ from Lemma 9.33 is neither empty (as is not abelian) nor equal to $A$ (as $a\not\in B$). Theorem 5.7 shows that $B$ 2-absorbs , which is a contradiction.
@@
Let $\mathbf{A}$ be a minimal Taylor algebra and $f_{1},...,f_{n}$ any collection of term operations of $\mathbf{A}$ such that for each edge $(a,b)$ with witnessing maximal congruence $\theta$ of $\operatorname{Sg}(a,b)$, at least one $f_{i}$ acts nontrivially on $\operatorname{Sg}(a,b)/\theta$ (that is, does not act as a projection). Then $f_{1},...,f_{n}$ generate the clone of $\mathbf{A}$.
Proof 9.47.
Let be the reduct of $\mathbf{A}$ with basic operations $f_{1},...,f_{n}$. By Proposition 2.2 it is enough to show that for any $a,b\in G$, there is no two-element quotient $\operatorname{Sg}(a,b)/\alpha$ such that each $f_{i}$ acts as a projection. Suppose for contradiction that there was such a pair $a,b$ and congruence $\alpha$ on $\operatorname{Sg}(a,b)$, and choose $a,b$ such that $\operatorname{Sg}(a,b)$ is inclusion minimal. Note that $a/\alpha$, $b/\alpha$, and $\operatorname{Sg}(a,b)$ might not be subuniverses of $\mathbf{A}$.
@@
If there is a congruence $\theta$ such that $\operatorname{Sg}_{\mathbf{A}}(a,b)/\theta$ is affine of prime order, then by assumption some $f_{i}$ acts nontrivially on $\operatorname{Sg}_{\mathbf{A}}(a,b)/\theta$, so there is some ternary term operation $p\in\operatorname{Clo}_{3}($ such that the restriction of $p$ to $\operatorname{Sg}_{\mathbf{A}}(a,b)/\theta$ is Mal’cev (by the fact that the affine Mal’cev algebra of $\mathbb{Z}/{p}$ has minimal clone which follows e.g. from the description of term operations of this algebra). Since $f_{i}$ acts as a projection on $\operatorname{Sg}(a,b)/\alpha$, $p$ also acts as a projection on $\operatorname{Sg}(a,b)/\alpha$. Suppose without loss of generality that $p$ does not act like third projection on $\operatorname{Sg}(a,b)/\alpha$. Then we have
@@
But then $\operatorname{Sg}(p(a,a,b),b)/\alpha=\operatorname{Sg}(a,b)/\alpha$ and
$\operatorname{Sg}(p(a,a,b),b)\subseteq\operatorname{Sg}(a,b)\cap b/\theta,$
contradicting the minimality of $\operatorname{Sg}(a,b)$ (as the latter set does not contain $a$ because $\theta$ separates $a$ and $b$).
Now suppose that $a$ is contained in a proper 3-absorbing subalgebra of $\operatorname{Sg}_{\mathbf{A}}(a,b)$. We show that this absorption is witnessed by an operation in $\operatorname{Clo}($.
If $C$ is a 2-absorbing subuniverse of $\operatorname{Sg}_{\mathbf{A}}(a,b)$, then there is some s-edge going into $C$ (by Proposition 9.22), and then some $f_{i}$ acts nontrivially on this edge, so there is a binary term operation $s$ of which acts as a semilattice operation on this edge. Since both coordinates of $s$ are essential and $C$ is a strongly projective subuniverse of $\operatorname{Sg}_{\mathbf{A}}(a,b)$ by Theorem 5.7, this operation witnesses the 2-absorption $C\trianglelefteq_{2}\operatorname{Sg}_{\mathbf{A}}(a,b)$.
If $C$ is not a 2-absorbing subuniverse of $\operatorname{Sg}_{\mathbf{A}}(a,b)$, then $C$ is not stable under majority edges by Lemma 9.25 and Theorem 5.19. Let $(c,d)$ with witnessing congruence $\beta$ be a majority edge such that $c\in C$ and $d/\beta\cap C=\emptyset$. Some $f_{i}$ acts nontrivially on $\operatorname{Sg}_{\mathbf{A}}(c,d)/\beta$, so there is a ternary term operation $m$ of that acts as the majority operation on $\operatorname{Sg}_{\mathbf{A}}(c,d)/\beta$. We claim that $m$ witnesses $C\trianglelefteq_{3}\operatorname{Sg}_{\mathbf{A}}(a,b)$. Indeed, $m\rightsquigarrow_{C}g$ for some $g\in\operatorname{Clo}(\{0,1\},\operatorname{maj})$ by Theorem 5.10, but $g$ cannot be a projection: e.g., if $g$ is the first projection, then $m(c,d,d)\in C$ (because $m$ is $\rightsquigarrow_{C}$-related to the first projection) and $m(c,d,d)\in d/\beta$ (because $m$ acts as the majority operation on the majority edge), a contradiction. Therefore $g=\operatorname{maj}$ and $m\rightsquigarrow_{C}\operatorname{maj}$ implies the claim.
In both cases, we see that there is some ternary operation $t\in\operatorname{Clo}($ which witnesses the 3-absorption $C\trianglelefteq_{3}\operatorname{Sg}_{\mathbf{A}}(a,b)$, and we may suppose without loss of generality that $t$ acts like the first projection on $\operatorname{Sg}(a,b)/\alpha$. Then we have
@@
But then $\operatorname{Sg}(a,t(b,a,a))/\alpha=\operatorname{Sg}(a,b)/\alpha$ and
$\operatorname{Sg}(a,t(b,a,a))\subseteq\operatorname{Sg}(a,b)\cap C,$
contradicting the minimality of $\operatorname{Sg}(a,b)$ again, and concluding the proof.
@@
Proof 9.48.
This is a consequence of Theorem 9.46.